Lauren purchases 500 shares of Walsh Corp. at the beginning of the year when the market price of those shares was $77.50. During the year she receives a dividend of $2.80 per share and sells the stock at the end of the year for $82.00 per share. Calculate Lauren's total percentage return on this investment if she: (a) pays for the shares in-full with cash; and, (b) uses the 60% margin her broker offers. Ignore commissions, interest charges and taxes
Answer #1

Initial money = 500 *77.5 = 38750

Dividends received = 500 *2.8 = 1400

Ending value = 82* 500 = 41000

If she uses margin, initial investment = 38750 *60% = 23250

a) Return when she pays for the shares in-full with cash = (41000 + 1400 -38750) / 38750 = 9.42%

b) Return when uses the 60% margin her broker offers = ( 41000+ 1400 - 38750) / 23250 = 15.70%